# Inferring Outcome Means of Exponential Family Distributions Estimated by Deep Neural Networks

### v2

@@ A.1 Preliminaries
$\frac{1}{2}\Delta_{n}(\widehat{f}_{n},f_{0})-c^{\prime}\bigg{(}\frac{\log\mathcal{N}_{n}}{n}+\bigg{(}\sqrt{\frac{\log\mathcal{N}_{n}}{n}}+1\bigg{)}\widetilde{\delta}\bigg{)}\quad\leq R_{n}(\widehat{f}_{n},f_{0})\leq 2\Delta_{n}(\widehat{f}_{n},f_{0})$
$+2\inf_{f\in\mathcal{F}(L,\mathbf{p},s,F)}\mathbb{E}\ell(\bm{X};f,f_{0})+C^{\prime}\bigg{(}\frac{\log\mathcal{N}_{n}}{n}+\bigg{(}\sqrt{\frac{\log\mathcal{N}_{n}}{n}}+1\bigg{)}\widetilde{\delta}\bigg{)}.$
The proof is provided in the Supplementary Material. To obtain the bounds in Proposition A.1, we apply truncation techniques that account for heterogeneity across individuals, paving the way for establishing convergence results under generalized nonparametric regression models. The following two lemmas provide properties of the covering number and characterize the minimal distance between functions $f\in\mathcal{F}(L,\mathbf{p},s,F)$ and $f_{0}$.
Lemma A.2.
@@
Proofs of Lemmas A.2 and A.3.
See Lemma 5 and the proof of Theorem 1 in Schmidt-Hieber (2020) for the proofs of Lemmas A.2 and A.3. ∎
@@
A.2 Proof of Theorem 3.5
With Lemma A.2 and if we let $\widetilde{\delta}=n^{-1}$, it holds that
@@
$\log(\mathcal{N}_{n})\precsim n\phi_{n}L\log^{2}n.$ (A.1)
By Proposition A.1, the lower bound for $R_{n}(\widehat{f}_{n},f_{0})$ can be written as
$\frac{1}{2}\Delta_{n}(\widehat{f}_{n},f_{0})-R_{n}(\widehat{f}_{n},f_{0})$ $\leq c^{\prime}\bigg{(}\frac{\log\mathcal{N}_{n}}{n}+\bigg{(}\sqrt{\frac{\log\mathcal{N}_{n}}{n}}+1\bigg{)}n^{-1}\bigg{)}$
@@
On the other hand, for the upper bound of $R_{n}(\widehat{f}_{n},f_{0})$, Lemma A.3 yields
@@
From Lemma A.4, it follows that $\mathbb{E}\ell(\bm{X};f,f_{0})\precsim\|f-f_{0}\|_{\infty}^{2}$. Hence, combining the upper bound of $\log(\mathcal{N}_{n})$ in (A.1) and $\inf_{f\in\mathcal{F}(L,\mathbf{p},s,F)}\|f-f_{0}\|_{\infty}^{2}$ above, we have
$R_{n}(\widehat{f}_{n},f_{0})-2\Delta_{n}(\widehat{f}_{n},f_{0})\precsim\phi_{n}L\log^{2}n.$
A.3 Proof of Corollary 3.6
According to Theorem 3.5, we have that when $\Delta_{n}(\widehat{f}^{b_{j}},f_{0})\leq C\phi_{r}L\log^{2}r$,
@@
Lemma A.4 further indicates that
@@
A.4 Proof of Theorem 3.7
From the assumption of
@@
$\mathcal{A}_{\delta}=\bigg{\{}\bm{X}\in\mathcal{X}:[\mathbb{E}\widehat{f}(\bm{X})-f_{0}(\bm{X})|\bm{X}]^{2}\leq\frac{C\phi_{r}L\log^{2}r}{\delta_{n}}\bigg{\}},$
@@
$\sqrt{\frac{n}{r^{2}\xi_{1,r}(\mathbf{x}_{*})}}\big{|}\mathbb{E}(\widehat{f}^{B}(\mathbf{x}_{*})-f_{0}(\mathbf{x}_{*}))\big{|}=O\bigg{(}\sqrt{\frac{n\phi_{r}L\log^{2}r}{\delta_{n}r^{2}\xi_{1,r}(\mathbf{x}_{*})}}\bigg{)}\to 0$ (A.2)
@@
Let $W_{b_{j}}$ be the number of subsamples $\mathcal{I}^{b_{j}}$ that contain index $i$, and define $g_{1,r}(x;\mathbf{x}_{*})=\mathbb{E}\widehat{f}^{b}(\mathbf{x}_{*};x,\mathbf{z}_{2},...,\mathbf{z}_{r})-\mathbb{E}\widehat{f}^{b}(\mathbf{x}_{*})$. We have $W_{b_{j}}$ is independent of $\mathbf{z}_{i}$ and $\mathbb{E}W_{b_{j}}=B{{n-1}\choose{r-1}}/{{n}\choose{r}}$. Hence, the equation above involving the term $\mathbf{z}_{i}$ can be further expressed as
@@
$\mathring{f}^{B}(\mathbf{x}_{*})=\frac{r}{n}\sum_{i=1}^{n}g_{1,r}(\mathbf{z}_{i};\mathbf{x}_{*}).$ (A.3)
Equation (A.3) provides a nice form, as $g_{1,r}(\mathbf{z}_{i})$ is independent of each other. Moreover, by the assumption that $\widehat{f}^{b}$ is bounded, the Lindeberg condition is satisfied, and it is clear that
$\frac{\sqrt{n}}{r\cdot\sqrt{\operatorname{{\rm Var}}(g_{1,r}(\mathbf{z}_{i};\mathbf{x}_{*}))}}\cdot\mathring{f}^{B}(\mathbf{x}_{*})~{}{\stackrel{{\scriptstyle d}}{{\rightarrow}}}~{}\mathcal{N}(0,1).$
By the definition of $\xi_{1,r}(\mathbf{x}_{*})$, it is easy to see that $\operatorname{{\rm Var}}(g_{1,r}(\mathbf{z}_{i};\mathbf{x}_{*}))=\xi_{1,r}(\mathbf{x}_{*})$, hence we have
$\sqrt{\frac{n}{r^{2}\xi_{1,r}(\mathbf{x}_{*})}}\cdot\mathring{f}^{B}(\mathbf{x}_{*})~{}{\stackrel{{\scriptstyle d}}{{\rightarrow}}}~{}\mathcal{N}(0,1).$ (A.4)
@@
$\widehat{f}^{b}(\mathbf{x}_{*};\mathbf{z}_{1},...,\mathbf{z}_{r})-\mathbb{E}\widehat{f}^{b}(\mathbf{x}_{*})=\sum_{i=1}^{r}T_{1}(\mathbf{z}_{i})+\sum_{i<j}T_{2}(\mathbf{z}_{i},\mathbf{z}_{j})...+T_{r}(\mathbf{z}_{1},...,\mathbf{z}_{r}).$ (A.5)
The $12^{r}-$ 1 random variables on the right hand side are all mean-zero and uncorrelated. Here, $T_{1}(\mathbf{z}_{i})=\mathbb{E}\big{(}\widehat{f}^{b}(\mathbf{x}_{*};\mathbf{z}_{1},...,\mathbf{z}_{r})|\mathbf{z}_{i}\big{)}-\mathbb{E}\widehat{f}^{b}(\mathbf{x}_{*})$. Applying the decomposition to $\widehat{f}^{B}(\mathbf{x}_{*})$ and utilizing the same trick we obtain (A.3), we have
$\mathbb{E}\big{(}\widehat{f}^{B}(\mathbf{x}_{*})-\mathbb{E}\widehat{f}^{b}(\mathbf{x}_{*})|\mathbf{z}_{1},...,\mathbf{z}_{n}\big{)}$ $=\frac{1}{{{n}\choose{r}}}\Bigg{(}{{n-1}\choose{r-1}}\sum_{i=1}^{n}T_{1}(\mathbf{z}_{i})+{{n-2}\choose{r-2}}\sum_{i<j}T_{2}(\mathbf{z}_{i},\mathbf{z}_{j})$
@@
$=\frac{1}{{{n}\choose{r}}}\Bigg{(}{{n-2}\choose{r-2}}\sum_{i<j}T_{2}(\mathbf{z}_{i},\mathbf{z}_{j})+...+\sum_{i_{1}<...<i_{r}}T_{r}(\mathbf{z}_{i_{1}},...,\mathbf{z}_{i_{r}})\Bigg{)}.$
Denote by $V_{k}=\operatorname{{\rm Var}}(T_{k}(\mathbf{z}_{i_{1}},...,\mathbf{z}_{i_{k}}))$, we then calculate the value of $\mathbb{E}\big{(}\widehat{f}^{B}(\mathbf{x}_{*})-\mathbb{E}\widehat{f}^{b}(\mathbf{x}_{*})-\mathring{f}^{B}(\mathbf{x}_{*})\big{)}^{2}$, which we will show later this term is negligible. To calculate this term, we apply the equality $\mathbb{E}X_{n}^{2}=\mathbb{E}\operatorname{{\rm Var}}(X_{n}|Y_{n})+\mathbb{E}(\mathbb{E}X_{n}|Y_{n})^{2}$. Then the term $\mathbb{E}\big{(}\widehat{f}^{B}(\mathbf{x}_{*})-\mathbb{E}\widehat{f}^{b}(\mathbf{x}_{*})-\mathring{f}^{B}(\mathbf{x}_{*})\big{)}^{2}$ can be written as
$\mathbb{E}\big{(}\widehat{f}^{B}(\mathbf{x}_{*})-\mathbb{E}\widehat{f}^{b}(\mathbf{x}_{*})-\mathring{f}^{B}(\mathbf{x}_{*})\big{)}^{2}$ $=\underbrace{\mathbb{E}\big{(}\operatorname{{\rm Var}}\big{(}\widehat{f}^{B}(\mathbf{x}_{*})-\mathbb{E}\widehat{f}^{b}(\mathbf{x}_{*})-\mathring{f}^{B}(\mathbf{x}_{*})|\mathbf{z}_{1},...,\mathbf{z}_{n}\big{)}\big{)}}_{J_{1}}$
@@
$J_{2}$ $=\frac{1}{{{n}\choose{r}}^{2}}\cdot\sum_{k=2}^{r}{{n-k}\choose{r-k}}^{2}\cdot{{n}\choose{k}}V_{k}$
$=\sum_{k=2}^{r}\frac{(n-k)!r!}{n!(r-k)!}{{r}\choose{k}}V_{k}$
$\leq\frac{r(r-1)}{n(n-1)}\cdot\sum_{k=2}^{r}{{r}\choose{k}}V_{k}$
$\leq\frac{r(r-1)}{n(n-1)}\operatorname{{\rm Var}}(\widehat{f}^{b}(\mathbf{x}_{*};\mathbf{z}_{1},...,\mathbf{z}_{r})).$
Here, the first equality is by the fact that all the terms are uncorrelated, the first inequality is by the decreasing monotonicity of $\frac{(n-k)!r!}{n!(r-k)!}$ with respect to $k$, and the last inequality is by the decomposition of $\widehat{f}^{b}$ in (A.5). To calculate $\mathbb{E}\big{(}\widehat{f}^{B}(\mathbf{x}_{*})-\mathbb{E}\widehat{f}^{b}(\mathbf{x}_{*})-\mathring{f}^{B}(\mathbf{x}_{*})\big{)}^{2}$, it remains for us to calculate $J_{1}$. Let $W_{i_{1}i_{2}...i_{k}}$ ($k\leq r$) be the number of the $B$ subsamples that contain index $i_{1},...,i_{k}$ simultaneously. Then we have
@@
$J_{1}$ $=\frac{1}{B^{2}}\sum_{k=2}^{r}\operatorname{{\rm Var}}(W_{i_{1}...i_{k}}){{n}\choose{k}}V_{k}$
$\leq\frac{1}{B}\sum_{k=2}^{r}\frac{{{n-k}\choose{r-k}}}{{{n}\choose{r}}}{{n}\choose{k}}V_{k}$
$=\frac{1}{B}\sum_{k=2}^{r}{{r}\choose{k}}V_{k}$
$\leq\frac{\operatorname{{\rm Var}}(\widehat{f}^{b}(\mathbf{x}_{*};\mathbf{z}_{1},...,\mathbf{z}_{r}))}{B}.$
Here, the first inequality is by $\operatorname{{\rm Var}}(W_{i_{1}...i_{k}})\leq B\cdot\frac{{{n-k}\choose{r-k}}}{{{n}\choose{r}}}$ and the uncorrelation on different terms $T_{k}$, and the last inequality is by $\sum_{k=2}^{r}{{r}\choose{k}}V_{k}\leq\operatorname{{\rm Var}}(\widehat{f}^{b}(\mathbf{x}_{*};\mathbf{z}_{1},...,\mathbf{z}_{r}))$. We hence conclude that
@@
$\leq\bigg{(}\frac{r(r-1)}{n(n-1)}+\frac{1}{B}\bigg{)}\operatorname{{\rm Var}}(\widehat{f}^{b}(\mathbf{x}_{*};\mathbf{z}_{1},...,\mathbf{z}_{r})),$
@@
$\frac{n}{r^{2}\xi_{1,r}(\mathbf{x}_{*})}\cdot\mathbb{E}\big{(}\widehat{f}^{B}(\mathbf{x}_{*})-\mathbb{E}\widehat{f}^{b}(\mathbf{x}_{*})-\mathring{f}^{B}(\mathbf{x}_{*})\big{)}^{2}$ $\leq\bigg{(}\frac{1}{n\xi_{1,r}(\mathbf{x}_{*})}+\frac{n}{Br^{2}\xi_{1,r}(\mathbf{x}_{*})}\bigg{)}\operatorname{{\rm Var}}(\widehat{f}^{b}(\mathbf{x}_{*};\mathbf{z}_{1},...,\mathbf{z}_{r}))$
$~{}{\stackrel{{\scriptstyle p}}{{\rightarrow}}}~{}0.$ (A.6)
Here, we utilize the fact that $\widehat{f}^{b}(\mathbf{x}_{*})$ is bounded. The bound of $\widehat{f}^{b}(\mathbf{x}_{*})$ indicates that $\operatorname{{\rm Var}}(\widehat{f}^{b}(\mathbf{x}_{*};\mathbf{z}_{1},\ldots,\mathbf{z}_{r}))$ is bounded. Additionally, we rely on the assumption that $\liminf_{n\to+\infty}n\xi_{1,r}(\mathbf{x}_{*})\to+\infty$ and $\liminf_{n\to+\infty}Br^{2}\xi_{1,r}(\mathbf{x}_{*})/n\to+\infty$. From (A.6), we have
@@
From (A.4), we also have that
@@
for $\mathbf{x}_{*}\in\mathcal{A}_{\delta_{n}}$. This is proved by (A.2) and $\mathbb{E}\widehat{f}^{B}(\mathbf{x}_{*})=\mathbb{E}\widehat{f}^{b}(\mathbf{x}_{*})$. Wrapping all together, we complete the proof of Theorem 3.7.
A.5 Proof of Theorem 3.8
Recall the definition of
$\widehat{V}_{i}=\frac{\sum_{j=1}^{B}(J_{b_{j}i}-J_{\cdot i})(\widehat{f}^{b_{j}}(\mathbf{x}_{*})-\widehat{f}^{B}(\mathbf{x}_{*}))}{B},\quad\widehat{\sigma}_{*}^{2}=\frac{n(n-1)}{(n-r)^{2}}\sum_{i=1}^{n}\widehat{V}_{i}^{2}.$
Without loss of generality, we calculate the term $\widehat{V}_{1}$, and the terms $\widehat{V}_{i}$ with $i\neq 1$ are exactly the same. Consider $\widehat{V}_{1}^{2}$ as $B/n$ tends to infinity. Then by the same combinatorics in the proof of Lemma 13 of Wager and Athey (2018), we can define

### v3

@@ C.1 Preliminaries
$\frac{1}{2}\Delta_{n}(\widehat{f}_{n})-c^{\prime}\bigg(\frac{\log\mathcal{N}_{n}}{n}+\bigg(\sqrt{\frac{\log\mathcal{N}_{n}}{n}}+1\bigg)\widetilde{\delta}\bigg)\quad\leq R_{n}(\widehat{f}_{n},f_{0})\leq 2\Delta_{n}(\widehat{f}_{n})$
[…]
The proof is provided in Section D. To obtain the bounds in Proposition C.1, we apply truncation techniques that account for heterogeneity across individuals, paving the way for establishing convergence results under generalized nonparametric regression models. The following two lemmas provide properties of the covering number and characterize the minimal distance between functions $f\in\mathcal{F}(L,\mathbf{p},s,F)$ and $f_{0}$.
Lemma C.2.
@@
Proofs of Lemmas C.2 and C.3.
See Lemma 5 and the proof of Theorem 1 in Schmidt-Hieber (2020) for the proofs of Lemmas C.2 and C.3. ∎
@@
C.2 Proof of Theorem 3.5
With Lemma C.2 and if we let $\widetilde{\delta}=n^{-1}$, it holds that
@@
$\log(\mathcal{N}_{n})\precsim n\phi_{n}L\log^{2}n.$ (C.1)
By Proposition C.1, the lower bound for $R_{n}(\widehat{f}_{n},f_{0})$ can be written as
$\frac{1}{2}\Delta_{n}(\widehat{f}_{n})-R_{n}(\widehat{f}_{n},f_{0})$ $\leq c^{\prime}\bigg(\frac{\log\mathcal{N}_{n}}{n}+\bigg(\sqrt{\frac{\log\mathcal{N}_{n}}{n}}+1\bigg)n^{-1}\bigg)$
@@
On the other hand, for the upper bound of $R_{n}(\widehat{f}_{n},f_{0})$, Lemma C.3 yields
@@
From Lemma C.4, it follows that $\mathbb{E}\ell(\bm{X};f,f_{0})\precsim\|f-f_{0}\|_{\infty}^{2}$. Hence, combining the upper bound of $\log(\mathcal{N}_{n})$ in (C.1) and $\inf_{f\in\mathcal{F}(L,\mathbf{p},s,F)}\|f-f_{0}\|_{\infty}^{2}$ above, we have
$R_{n}(\widehat{f}_{n},f_{0})-2\Delta_{n}(\widehat{f}_{n})\precsim\phi_{n}L\log^{2}n.$
C.3 Proof of Corollary 3.6
According to Theorem 3.5, we have that when $\Delta_{n}(\widehat{f}^{b_{j}})\leq C\phi_{r}L\log^{2}r$,
@@
Lemma C.4 further indicates that
@@
C.4 Proof of Theorem 3.7
By Assumption 3.4, recall $\gamma>\frac{1}{1+\min_{j=0,...,q}(2\gamma_{j}^{*})/(2\gamma_{j}^{*}+t_{j})}$. For the sufficiently small $\varepsilon>0$, it is easy to see that
@@
$\mathcal{A}_{\delta_{n}}=\bigg\{\bm{X}\in\mathcal{X}:[\mathbb{E}\widehat{f}(\bm{X})-f_{0}(\bm{X})|\bm{X}]^{2}\leq\frac{C\phi_{r}L\log^{2}r}{\delta_{n}}\bigg\},$
@@
$\sqrt{\frac{n}{r^{2}\xi_{1,r}(\mathbf{x}_{*})}}\big|\mathbb{E}(\widehat{f}^{B}(\mathbf{x}_{*})-f_{0}(\mathbf{x}_{*}))\big|=O\bigg(\sqrt{\frac{n\phi_{r}L\log^{2}r}{\delta_{n}r^{2}\xi_{1,r}(\mathbf{x}_{*})}}\bigg)\to 0$ (C.2)
@@
Let $W_{b_{j}}$ be the number of subsamples $\mathcal{I}^{b_{j}}$ that contain index $i$, and define $g_{1,r}(\mathbf{z};\mathbf{x}_{*})=\mathbb{E}\widehat{f}^{b}(\mathbf{x}_{*};\mathbf{z},\mathbf{z}_{2},...,\mathbf{z}_{r})-\mathbb{E}\widehat{f}^{b}(\mathbf{x}_{*})$. We have $W_{b_{j}}$ is independent of $\mathbf{z}_{i}$ and $\mathbb{E}W_{b_{j}}=B{\binom{n-1}{r-1}}/{\binom{n}{r}}$. Hence, the equation above involving the term $\mathbf{z}_{i}$ can be further expressed as
@@
$\mathring{f}^{B}(\mathbf{x}_{*})=\frac{r}{n}\sum_{i=1}^{n}g_{1,r}(\mathbf{z}_{i};\mathbf{x}_{*}).$ (C.3)
Equation (C.3) provides a nice form, as $g_{1,r}(\mathbf{z}_{i})$ is independent of each other. Moreover, by the assumption that $\widehat{f}^{b}$ is bounded, the Lindeberg condition is satisfied, and it is clear that
$\frac{\sqrt{n}}{r\cdot\sqrt{\mathrm{Var}(g_{1,r}(\mathbf{z}_{i};\mathbf{x}_{*}))}}\cdot\mathring{f}^{B}(\mathbf{x}_{*})~{\stackrel{{\scriptstyle d}}{{\rightarrow}}}~\mathcal{N}(0,1).$
By the definition of $\xi_{1,r}(\mathbf{x}_{*})$, it is easy to see that $\mathrm{Var}(g_{1,r}(\mathbf{z}_{i};\mathbf{x}_{*}))=\xi_{1,r}(\mathbf{x}_{*})$, hence we have
$\sqrt{\frac{n}{r^{2}\xi_{1,r}(\mathbf{x}_{*})}}\cdot\mathring{f}^{B}(\mathbf{x}_{*})~{\stackrel{{\scriptstyle d}}{{\rightarrow}}}~\mathcal{N}(0,1).$ (C.4)
@@
$\widehat{f}^{b}(\mathbf{x}_{*};\mathbf{z}_{1},...,\mathbf{z}_{r})-\mathbb{E}\widehat{f}^{b}(\mathbf{x}_{*})=\sum_{i=1}^{r}T_{1}(\mathbf{z}_{i})+\sum_{i<j}T_{2}(\mathbf{z}_{i},\mathbf{z}_{j})...+T_{r}(\mathbf{z}_{1},...,\mathbf{z}_{r}).$ (C.5)
The $2^{r}-1$ random variables on the right hand side are all mean-zero and uncorrelated. Here, $T_{1}(\mathbf{z}_{i})=\mathbb{E}\big(\widehat{f}^{b}(\mathbf{x}_{*};\mathbf{z}_{1},...,\mathbf{z}_{r})|\mathbf{z}_{i}\big)-\mathbb{E}\widehat{f}^{b}(\mathbf{x}_{*})$. Applying the decomposition to $\widehat{f}^{B}(\mathbf{x}_{*})$ and utilizing the same trick we obtain (C.3), we have
$\mathbb{E}\big(\widehat{f}^{B}(\mathbf{x}_{*})-\mathbb{E}\widehat{f}^{b}(\mathbf{x}_{*})|\mathbf{z}_{1},...,\mathbf{z}_{n}\big)$ $=\frac{1}{{\binom{n}{r}}}\Bigg({\binom{n-1}{r-1}}\sum_{i=1}^{n}T_{1}(\mathbf{z}_{i})+{\binom{n-2}{r-2}}\sum_{i<j}T_{2}(\mathbf{z}_{i},\mathbf{z}_{j})$
@@
$=\frac{1}{{\binom{n}{r}}}\Bigg({\binom{n-2}{r-2}}\sum_{i<j}T_{2}(\mathbf{z}_{i},\mathbf{z}_{j})+...+\sum_{i_{1}<...<i_{r}}T_{r}(\mathbf{z}_{i_{1}},...,\mathbf{z}_{i_{r}})\Bigg).$
Denote by $V_{k}=\mathrm{Var}(T_{k}(\mathbf{z}_{i_{1}},...,\mathbf{z}_{i_{k}}))$, we then calculate the value of $\mathbb{E}\big(\widehat{f}^{B}(\mathbf{x}_{*})-\mathbb{E}\widehat{f}^{b}(\mathbf{x}_{*})-\mathring{f}^{B}(\mathbf{x}_{*})\big)^{2}$, which we will show later this term is negligible. To calculate this term, we apply the equality $\mathbb{E}X_{n}^{2}=\mathbb{E}\mathrm{Var}(X_{n}|Y_{n})+\mathbb{E}(\mathbb{E}X_{n}|Y_{n})^{2}$. Then the term $\mathbb{E}\big(\widehat{f}^{B}(\mathbf{x}_{*})-\mathbb{E}\widehat{f}^{b}(\mathbf{x}_{*})-\mathring{f}^{B}(\mathbf{x}_{*})\big)^{2}$ can be written as
$\mathbb{E}\big(\widehat{f}^{B}(\mathbf{x}_{*})-\mathbb{E}\widehat{f}^{b}(\mathbf{x}_{*})-\mathring{f}^{B}(\mathbf{x}_{*})\big)^{2}$ $=\underbrace{\mathbb{E}\big(\mathrm{Var}\big(\widehat{f}^{B}(\mathbf{x}_{*})-\mathbb{E}\widehat{f}^{b}(\mathbf{x}_{*})-\mathring{f}^{B}(\mathbf{x}_{*})|\mathbf{z}_{1},...,\mathbf{z}_{n}\big)\big)}_{J_{1}}$
@@
$J_{2}$ $=\frac{1}{{\binom{n}{r}}^{2}}\cdot\sum_{k=2}^{r}{\binom{n-k}{r-k}}^{2}\cdot{\binom{n}{k}}V_{k}$
$=\sum_{k=2}^{r}\frac{(n-k)!r!}{n!(r-k)!}{\binom{r}{k}}V_{k}$
$\leq\frac{r(r-1)}{n(n-1)}\cdot\sum_{k=2}^{r}{\binom{r}{k}}V_{k}$
$\leq\frac{r(r-1)}{n(n-1)}\mathrm{Var}(\widehat{f}^{b}(\mathbf{x}_{*};\mathbf{z}_{1},...,\mathbf{z}_{r})).$
Here, the first equality is by the fact that all the terms are uncorrelated, the first inequality is by the decreasing monotonicity of $\frac{(n-k)!r!}{n!(r-k)!}$ with respect to $k$, and the last inequality is by the decomposition of $\widehat{f}^{b}$ in (C.5). To calculate $\mathbb{E}\big(\widehat{f}^{B}(\mathbf{x}_{*})-\mathbb{E}\widehat{f}^{b}(\mathbf{x}_{*})-\mathring{f}^{B}(\mathbf{x}_{*})\big)^{2}$, it remains for us to calculate $J_{1}$. Let $W_{i_{1}i_{2}...i_{k}}$ ($k\leq r$) be the number of the $B$ subsamples that contain index $i_{1},...,i_{k}$ simultaneously. Then we have
@@
$J_{1}$ $=\frac{1}{B^{2}}\sum_{k=2}^{r}\mathrm{Var}(W_{i_{1}...i_{k}}){\binom{n}{k}}V_{k}$
$\leq\frac{1}{B}\sum_{k=2}^{r}\frac{{\binom{n-k}{r-k}}}{{\binom{n}{r}}}{\binom{n}{k}}V_{k}$
$=\frac{1}{B}\sum_{k=2}^{r}{\binom{r}{k}}V_{k}$
$\leq\frac{\mathrm{Var}(\widehat{f}^{b}(\mathbf{x}_{*};\mathbf{z}_{1},...,\mathbf{z}_{r}))}{B}.$
Here, the first inequality is by $\mathrm{Var}(W_{i_{1}...i_{k}})\leq B\cdot\frac{{\binom{n-k}{r-k}}}{{\binom{n}{r}}}$ and the uncorrelation on different terms $T_{k}$, and the last inequality is by $\sum_{k=2}^{r}{\binom{r}{k}}V_{k}\leq\mathrm{Var}(\widehat{f}^{b}(\mathbf{x}_{*};\mathbf{z}_{1},...,\mathbf{z}_{r}))$. We hence conclude that
@@
$\leq\bigg(\frac{r(r-1)}{n(n-1)}+\frac{1}{B}\bigg)\mathrm{Var}(\widehat{f}^{b}(\mathbf{x}_{*};\mathbf{z}_{1},...,\mathbf{z}_{r})),$
@@
$\frac{n}{r^{2}\xi_{1,r}(\mathbf{x}_{*})}\cdot\mathbb{E}\big(\widehat{f}^{B}(\mathbf{x}_{*})-\mathbb{E}\widehat{f}^{b}(\mathbf{x}_{*})-\mathring{f}^{B}(\mathbf{x}_{*})\big)^{2}$ $\leq\bigg(\frac{1}{n\xi_{1,r}(\mathbf{x}_{*})}+\frac{n}{Br^{2}\xi_{1,r}(\mathbf{x}_{*})}\bigg)\mathrm{Var}(\widehat{f}^{b}(\mathbf{x}_{*};\mathbf{z}_{1},...,\mathbf{z}_{r}))$
$~{\stackrel{{\scriptstyle p}}{{\rightarrow}}}~0.$ (C.6)
Here, we utilize the fact that $\widehat{f}^{b}(\mathbf{x}_{*})$ is bounded. The bound of $\widehat{f}^{b}(\mathbf{x}_{*})$ indicates that $\mathrm{Var}(\widehat{f}^{b}(\mathbf{x}_{*};\mathbf{z}_{1},\ldots,\mathbf{z}_{r}))$ is bounded. Additionally, we rely on Assumption 3.4 that $\liminf_{n\to+\infty}n\xi_{1,r}(\mathbf{x}_{*})\geq n^{1-\varepsilon-\gamma}\to+\infty$ and $\liminf_{n\to+\infty}Br^{2}\xi_{1,r}(\mathbf{x}_{*})/n\to+\infty$. From (C.6), we have
@@
From (C.4), we also have that
@@
for $\mathbf{x}_{*}\in\mathcal{A}_{\delta_{n}}$. This is proved by (C.2) and $\mathbb{E}\widehat{f}^{B}(\mathbf{x}_{*})=\mathbb{E}\widehat{f}^{b}(\mathbf{x}_{*})$. Wrapping all together, we complete the proof of Theorem 3.7.
C.5 Proof of Theorem 3.8
Recall the definition
$Z_{b_{j}i}=(J_{b_{j}i}-J_{\cdot i})(\widehat{f}^{b_{j}}(\mathbf{x}_{*})-\widehat{f}^{B}(\mathbf{x}_{*})),\quad\widehat{V}_{i}=\frac{\sum_{j=1}^{B}Z_{b_{j}i}}{B}.$
We denote by
$\widetilde{\sigma}^{2}=\frac{n(n-1)}{(n-r)^{2}}\sum_{i=1}^{n}\widehat{V}_{i}^{2}.$
Without loss of generality, we calculate the term $\widehat{V}_{1}$, and the terms $\widehat{V}_{i}$ with $i\neq 1$ are exactly the same. Consider $\widehat{V}_{1}^{2}$ and $B/n\to c$ for some constant $c>0$. Then by the inspiration of the proof of Lemma 13 of Wager & Athey (2018), we can define